1a. If the present value of $157 is $143, what is the discount factor? (Round your answer to 4 decimal places.)

b. If that $157 is received in year 5, what is the interest rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

2. Suppose you own a small company that is contemplating construction of a suburban office block. The cost of buying the land and constructing the building is $720,000. Your company has cash in the bank to finance construction. Your real estate adviser suggests that you rent out the building for two years at $31,000 a year and predicts that at the end of that time you will be able to sell the building for $848,000.

Thus there are now two future cash flows--a cash flow of C1 = $31,000 at the end of year 1 and a further cash flow of C2 = ($31,000 + 848,000) = $879,000 at the end of the second year.

a. Calculate the NPV of the office building venture at interest rates of 6, 11, and 16%. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 2 decimal places.)

Net Present value at 5%

Net Present value at 10%

Net Present value at 15%

b. At what discount rate (approximately) would the project have a zero NPV? Check your answer by calculating the NPV at your approximate rate; it should be close to zero. (Enter your answer as a percent rounded to the nearest whole number.)
Answer #1
1.a. Discount factor if PV of $157 is $143
=(157/143)-1 0.07692308
1.b. interest rate if $157 is received after 5 years
=(157/143)^(1/5)-1 0.01886
2.a. Computation of NPV of the office building
Y0 Y1 Y2 NPV
Discount rate @6% 1.000 0.943 0.890
Cash flows -720000 31000 879000
PV of Cash flows -720000 29245.28302 782306.8708 91552.15
Y0 Y1 Y2 NPV
Discount rate @11% 1.000 0.901 0.812
Cash flows -720000 31000 879000
PV of Cash flows -720000 27927.92793 713416.1188 21344.05
Y0 Y1 Y2 NPV
Discount rate @16% 1.000 0.862 0.743
Cash flows -720000 31000 879000
PV of Cash flows -720000 26724.13793 653240.1902 -40035.7
2.b. Discount rate at which NPV is 0 =(879000/(1+IRR)^2)+(31000/(1+IRR)) -720000 13%